Calculate the molar mass of the following substances.
Nitric acid, HNO3

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

Nitric acid, HNO3 = 1 × 1 + 1 × 14 + 3 × 16
= 1 + 14 + 48 = 63 g

Thanks

Shraddha Singh
2 years ago
Dear Student,
 
Molar mass of  HNO= Mass of H + Mass of N + Mass of O × 2
= 1 + 14 + (16×2)
= 15 + 32
= 47 g / mol
 
Hope it helps!
Thank you